OpenStudy (love.kat123):

NEED HELP ASAP How many solutions does the system of equations have? 2x + 3y = 6 y=-2/3x-1 A. 0 B. 1 C. 2 D. infinite

OpenStudy (texaschic101):

first make sure both equations are in y = mx + b form. Can you do that ?

OpenStudy (tkhunny):

Have you considered finding the solutions?

OpenStudy (love.kat123):

tbh i dont know how to do this at all

OpenStudy (tkhunny):

2x + 3y = 6 y=-2/3x-1 OR any other form, so long as they are consistent. One in Slope-Intercept and the other in Standard just isn't helpful. y = (-2/3)x - 1 3y = -2x - 1 2x + 3y = -1 This is now easily compared to 2x + 3y = 6

OpenStudy (tkhunny):

You MUST do better than that. If you do not know how to do it at all, why were you given this problem? This situation is very unlikely to be as you have described it. The idea is not to have the solution fall out of the sky. You must think it through. You must use the rules, properties, and techniques you have been taught.

OpenStudy (texaschic101):

In y = mx + b form, the number in the m position is the slope and the number in the b position is the y intercept. So first we will put both equations in y = mx + b form and compare the slopes and the y intercepts. y = -2/3x - 1 ---- already in y = mx + b form slope = -2/3 and y intercept = -1 2x + 3y = 6 3y = -2x + 6 y = -2/3x + 2 slope is -2/3 and y intercept is 2 now remember this ... if slopes are the same and y intercepts different....parallel line and no solution same slope and same y intercept...same line...infinite solutions different slope and different y intercepts means it has 1 solution so what do you think now ?

OpenStudy (love.kat123):

i think u guys r confusing

OpenStudy (texaschic101):

I first put the equations in y = mx + b form where m is your slope and b is your y intercept. You can then compare the slopes and the y intercepts. By following the rules I listed above and doing the comparisons, you should find that they have same slope and different y intercepts, thus, they are parallel and have no solutions.

OpenStudy (tkhunny):

Nah. Just me. texaschic101 is awesome. Really, we're just trying to show you one way of thinking that will solve the problem. We did it a little differently. That's fine. You need to find a way to think about it - a way you understand. You have two examples of possible ways to think about it.